### v1

# A comprehensive framework for statistical testing of brain dynamics

Abstract
We introduce a comprehensive statistical framework for analysing brain dynamics and testing their associations with behavioural, physiological and other non-imaging variables. Based on a generalisation of the Hidden Markov Model (HMM) — the Gaussian-Linear HMM — our open-source Python package supports multiple experimental paradigms, including task-based and resting-state studies, and addresses a wide range of questions in neuroscience and related scientific fields. Inference is carried out using permutation-based methods and structured Monte Carlo resampling, and the framework can easily handle confounding variables, multiple testing corrections, and hierarchical relationships within the data. The package includes tools for intuitive visualisation of statistical results, along with comprehensive documentation and step-by-step tutorials to make it accessible for users of varying expertise. Altogether, it provides a broadly applicable, end-to-end pipeline for analysis and statistical testing of functional neural data and its dynamics.
Keywords: Brain dynamics, statistical inference, permutation testing, neuroscience, Hidden Markov model, neuroimaging, electrophysiology
Introduction
A fundamental aspect of neuroscience is the necessity to find, assess, and characterise associations between brain activity and behaviour[1, 2]. Understanding these brain-behaviour relationships is crucial for advancing our knowledge of cognitive functions, mental health, and neurological disorders[3, 4, 5]. To quantify such associations, researchers typically rely on either prediction techniques or statistical testing. While prediction methods focus on out-of-sample accuracy, explanatory approaches rely on statistical testing to determine whether specific patterns of interest, such as the relationship between a measurement obtained from the brain and a behavioural trait, are supported by the data[6]. Here, we present easy-to-use and well-documented routines for statistical testing between brain dynamics and different forms of behaviour. These tools can readily be applied in other domains.
The recently introduced Gaussian-Linear Hidden Markov Model (GLHMM) Python package was developed to characterise brain dynamics across multiple data modalities through the identification of latent brain states[7]. This package includes a versatile statistical testing framework that supports a wide range of experimental designs commonly used in neuroscience research, including the resting state, often available in public datasets. With an emphasis on statistical testing, this paper provides a detailed overview of the entire analysis protocol, from model fitting to result presentation.
We define four families of statistical tests that address a wide range of relevant questions in neuroscience. These are: (1) across-subjects tests, which assess associations between individual traits and brain activity across-subjects; (2) across-trials tests, which compare brain activity across trials under different experimental conditions; (3) across-session within-subject tests, which evaluate long-term changes in brain dynamics across multiple sessions for one subject; and (4) across-state-visits tests, which examine associations between brain time series and one or more simultaneously measured variables, such as physiological measurements.
Unlike existing frameworks for statistical inference that primarily target time-averaged or non-temporal data, the present protocol has a strong focus on the temporal dimension of brain activity (i.e. on brain dynamics), even though it can also handle more conventional tests. While these tests are presented with a focus in neuroscience, they can readily be adapted to other fields such as economics, ecology, and other fields. This makes the toolbox suitable for any domain that involves statistical testing of relationships between dynamic system properties (e.g., sequential or temporal data) and a set of external variables.
Development of the protocol
\@fb@secFB
Estimating a model brain dynamics from time series data
This protocol examines the relationship between brain dynamics and behaviour at various temporal scales through statistical testing. Brain dynamics are first characterised using the GLHMM, a generalisation of the Hidden Markov Model (HMM), before proceeding to the statistical testing, which forms the main focus of this paper. The HMM characterises brain activity using a finite set of latent states and their temporal dynamics, i.e. when they occur and the transitions between them. The GLHMM extends the standard Gaussian-state HMM by allowing multiple types of state models based on different configurations of the regression model. Leveraging this flexibility, it can be used on different brain modalities, including functional MRI (fMRI)[8, 9, 10, 11, 12], magnetoencephalography (MEG)[13, 14, 15, 16], electroencephalography (EEG)[17, 13], electrocorticography (ECoG)[7], and local field potentials (LFPs)[18, 19]. Specifically, the GLHMM is based on a Bayesian regression model to capture the relationship between two time series: $X$ (independent variable) and $Y$ (dependent variable). The observations are modelled as:
$Y_{t}|s_{t}=k\sim N(\mu_{k}+X_{t}\beta_{k},\Sigma_{k})$
where $s_{t}$ is a variable indicating which state is active at time point $t$, $\mu_{k}$ is the baseline activity for state $k$, $\beta_{k}$ represents the regression coefficients linking $X$ to $Y$ for state $k$, and $\Sigma_{k}$ is the covariance matrix for state $k$. This allows for flexible modelling of the data, where the parameters $\mu_{k}$, $\beta_{k}$, and $\Sigma_{k}$ may vary across states, remain global, or be not modelled (in the case of the covariance matrix, this corresponds to using the identity matrix). This model reduces to the standard Gaussian HMM when $\beta_{k}$ is unmodelled, and $\mu_{k}$ is state-specific. Furthermore, the transition probabilities describe the likelihood of switching from one state to another:
$P(s_{t}=k\,|\,s_{t-1}=l)$
To estimate the posterior distribution of the model parameters, including the state time courses, represented as the probabilities $\gamma_{tk}=P(s_{t}=k|s_{t-1}=l,X_{t},Y_{t})$, the GLHMM uses variational inference.
Overall, thanks to its flexible parametrisation, the GLHMM allows for many time-varying functional connectivity analyses, at the whole-brain level or targeting specific connections or networks[20].
\@fb@secFB
Performing statistical testing on the estimated model of brain dynamics
Once the model has been fitted to the data, we will use formal statistical testing to examine the associations between the model parameters (representing different aspects of the time series’ dynamics) and the behavioural or experimental variables. To assess whether these associations are statistically meaningful, the framework primarily relies on permutation-based inference, which does not impose any distributional assumption. This avoids issues when these assumptions are violated, which can lead to unreliable p-values and inflated false-positive rates. Alongside permutation-based methods, the framework also includes a test that use structured Monte Carlo resampling, the across-state-visits test, which is discussed below.
As represented in Figure 1, the four types of tests presented in this protocol are: across-subjects, across-trials-within-session, across-session-within-subject, and across-state-visits. We next succinctly describe the four tests, and further details can be found in the SI.
Figure 1: Illustration of the four statistical tests. (A) the across-subjects test compares behavioural measurements (or traits) across multiple individuals to test subject trait differences. (B) The across-trials test assesses differences in brain responses across experimental conditions, such as two types of stimuli. (C) The across-session-within-subject test assesses changes in brain responses over experimental sessions, given an experimental paradigm such as the one used in B. (D) The across-state-visits test assesses relationships between state time courses and concurrently recorded signals, where each state may correspond to the activation a specific brain network.
The across-subjects test considers data from multiple individuals (or brain scans) to assess the associations between subject-specific model parameters encoding different aspects of brain dynamics on the one hand, and one or more subject-specific non-imaging traits (e.g. age, sex, cognitive capacity) on the other hand. For instance, we might be interested in testing the relationship between the time spent in the default mode network state (DMN) at wakeful rest and a clinical trait such as anxiety levels, cognitive decline, or depression risk. For permutation testing, an important requirement is the exchangeability of subjects or scans, meaning that, after permutation, in the absence of real effect, the distribution of the data remains the same as that of the unpermuted. However, if subjects have familial relationships, this assumption would be violated, making the test invalid. We address this issue in two ways. In the simplest scenario, the subjects or scans can be assigned to blocks, such that permutations are carried out at the block level, either within or between blocks. For example, if we had several scans per subject and one non-imaging trait per scan, the blocks would correspond to the subjects. In more complex scenarios, nested relationships among subjects or different types of relationships can be considered. Here, the user provides a hierarchical tree to account for this structure in the permutation scheme; for more details, see [21].
The across-trials test considers experimental studies where subjects perform a task across multiple trials within a single session or multiple sessions to assess differences in the states’ time courses between experimental conditions (or subject actions). A typical example is a visual paradigm comparing two types of stimuli, where the goal is to identify when network activity significantly differs between conditions. The test generates a surrogate (null) distribution by performing permutations only on trials within the same session in which they were recorded. This test can be run at each time point throughout the trial to examine how the effects unfold over time.
The across-sessions within-subjects test provides a new way to assess whether the brain-behaviour relationship under study changes over slower time scales, i.e. over the course of multiple sessions. This approach can be used in longitudinal studies where a subject is scanned repeatedly while performing a task involving one or more contrasts (e.g. stimuli or subject decisions). Unlike traditional approaches that shuffle trial data, this method operates at the level of regression coefficients. For each session, a regression model is fitted to estimate session-specific beta coefficients, which capture the relationship between brain activity (here, state time courses) and the experimental condition. To test for significant changes across sessions, the method generates a (null) distribution by randomly permuting these beta coefficients across sessions (rather than permuting the data). This approach accounts for differences in session length, variations in condition proportions, and the lack of direct alignment between trials across sessions (see S1 for more details). Similar to the across-trials test, this can be performed at each time point to produce a time-resolved statistical analysis.
The across-state-visits test, also novel in this context, evaluates whether the state time courses (represented by the Viterbi path, a discretised version of the state activation probabilities $\gamma_{tk}$) are associated with concurrently recorded physiological or behavioural signals over time, such as pupil size, heart rate, or skin conductance. For instance, one state might correspond to an increase in pupil size, while another might correspond to a decrease. The Viterbi path thus serves as the contrast, enabling the comparison of differences in the second set of signals. Standard permutation methods are however not suitable for this test, as shuffling time points would disrupt the temporal structure of the data. Instead, using a Monte Carlo approach, we generate surrogate Viterbi paths that preserve the original transition timings but randomly reassign which states are visited at each transition in a structured manner. This is done such that it maintains the statistical properties of the original data while breaking the observed association between states and the external signal, with further elaboration on this test available in S2.
Together, these tests provide a framework for systematically studying dynamic brain-behavior relationships. The protocol below describes each step in detail, addressing the key challenges user can encounter.
Application of the method
The strength of the HMM framework comes from its ability to (i) detect fast changes in the properties of the data in a data-driven manner and (ii) offer a cohesive representation of dynamics at both group and subject level using a well-defined model. Leveraging this capacity, the HMM has been used to investigate several neuroscience questions in recent years, such as: the nature of the sleep cycle from fMRI data[10]; the long-term temporal structure of key brain networks during spontaneous cognition using MEG[8]; the relationship between the temporal patterns of whole-brain networks and the spontaneous replay of previously learned sequences in MEG[16]; the dynamics of memory retrieval throughout the cortex in fMRI[22]; the spectral characterisation of large-scale cortical networks at rest in MEG[15]; and how fast-changing brain states relate to specific social behaviour dynamics[23]. The comprehensive set of statistical tests introduced here can streamline and systematise the investigation of these and related questions, facilitating the exploration of associations between brain state dynamics and behaviour.
However, the methods in this protocol are sufficiently general that are not limited to neuroscience, and can be applied to other fields. For instance, in economics, these may be used to identify periods when key political events correlate more strongly with economic indices, such as inflation, unemployment rates, and gross domestic product (GDP). For example, during major political events like elections, shifts in government policies, or international trade agreements, economic indicators may show increased correlation as businesses and consumers react to potential changes in regulation, taxation, or trade relationships. Modelling these relationships can provide a better understanding of how such events influence consumer behaviour and economic dynamics. In ecology, as another example, this protocol could be used to study animal migration patterns by identifying latent states that correspond to different stages of migration, such as foraging, resting, or travelling. These states can be further analysed to understand how they change in response to environmental factors such as food availability. For instance, shifts in migration routes or timing may be linked to climate change or human activities.
From a practical point of view, this protocol supports a range of industry-standard data formats, including CSV, text files, and NIFTI, owing to Python’s robust data handling capabilities. Given this versatility, integrating data from other fields into the framework is straightforward and efficient. Documentation is available at ’Read the Docs’, including tutorials and examples.
In summary, the presented protocol has the potential to aid research in multiple fields involving temporal data by identifying latent factors underlying the dynamics of complex systems. The tests can also be applied generally to any time series, regardless of whether the HMM is used.
Experimental design
As mentioned above, the statistics toolbox of this protocol includes four tests: across-subjects, across-trials, across-sessions-within-subject, and across-state-visits. Figure 2 presents an overview of the procedure for applying these tests using the toolbox. The procedure is divided into three parts: preparing the data, applying the statistical analysis, and visualising the results. While part 1 includes steps 2–5 that are specific to the HMM model, the statistical tests themselves can be applied to any data type and do not require HMM outputs.
Figure 2: Schematic of the analysis pipeline. State time courses estimated by the HMM from brain data, which may be used as inputs for the statistical testing procedure. Part 1b) The input data used for statistical testing between matrices $D$ and $R$. The HMM output is here used to generate $D$; however, this is optional, and $D$ can represent any data with the appropriate structure as specified in the protocol. Matrix $R$ represents behavioural or other non-imaging data. Part 2) The statistical testing procedure produces a matrix of p-values, whose shape depends on the specific test applied. Part 3) Visualisation of the results using different types of plots.
\@fb@secFB
Install and set-up Python environment
Before beginning the analysis, we set up a Python environment and install the required packages. We start by creating a Python environment to manage dependencies separately from other projects. Once the environment is ready, we install the package by running the following command in the terminal:
pip install glhmm
This command clones the GitHub repository with all the code required for the procedure. The protocol can be run on a local computer or, if needed, on Google Colab.
\@fb@secFB
Load and prepare data: Step 1 - Load data into Python environment
To begin the analysis, we load brain and behavioural data. The aim is to examine potential associations between brain activity and behaviour using these two types of data. Functional brain data can come from different techniques, such as fMRI, EEG, MEG, LFP, or ECoG (although the statistical tests are general enough to be directly applied to structural measures as well). Behavioural data, or more generally, non-imaging data can be cognitive or demographic information, or any clinical variable. While the brain data can go through extra processing using the HMM model (described in steps 2-5), these steps are optional. That is, any imaging set of variables, even if not produced by the HMM, can be used as $D$ (see Figure 2).
\@fb@secFB
Load and prepare data: Step 2 - Data structuring for the HMM (optional)
When the brain data is prepared for training an HMM model, it needs to be shaped as a [(No. of timepoints $\boldsymbol{\cdot}$ No. of subjects/sessions) $\times$ No. of features] matrix. In this format, the data from all subjects or sessions are combined along the first dimension, while the second dimension represents the features, such as brain regions or channels. If the brain data is provided (e.g., [No. of timepoints, No. of subjects, No. of features]) tensor, we can reshape it by concatenating timepoints and subjects or sessions along one dimension, with the features remaining as the second dimension.
\@fb@secFB
Load and prepare data: Step 3 - Preprocessing data (optional)
Before analysis, the raw data may need to be cleaned to remove noise and artefacts. The package offers tools for basic preprocessing, such as standardising the data (to keep measurements on the same scale), filtering (for noise removal or to isolate specific frequency bands), and dimensionality reduction (using PCA or ICA). If additional preprocessing steps are needed, these should be handled separately.
\@fb@secFB
Load and prepare data: Step 4 - Set-up and train a HMM (optional)
The next steps are initialising and training the HMM with preprocessed data. Before training, the number of these states needs to be defined, based on the needs of the analysis and the size of the data[15, 12, 19]. Also, the type of state model has to be chosen [7]. Once trained, the model saves the learned parameters and state time courses (referred to as ’gamma’ in the code), which represent the probability of each state to occur at each time point. These state time course values are used in subsequent statistical tests to examine how state transitions relate to cognitive and behavioural measures.
\@fb@secFB
Load and prepare data: Step 5 - Configure HMM outputs for statistical analysis (optional)
The HMM output takes different forms, depending on the type of test we wish to carry out. By default, it produces continuous state time courses, which can be used to study changes over the full recording. Alternatively, the state time courses can be epoched, creating a 3D tensor, or summarised into a 2D matrix of aggregated statistics. When epoching is applied, the state time courses are divided into segments based on specific experimental events (i.e. trials), such as responses to stimuli or other time-locked occurrences. This allows for the analysis of how brain states differ during these targeted periods. Finally, we can compute some form of aggregated statistics to generate a single set of values for each subject or session, summarising the main patterns in the state time courses across the whole series or within specific time windows. These statistics include fractional occupancy (FO), which represents the proportion of time spent in each state during a given period; dwell time, the average duration spent in a state, reflecting its stability; switching rate, the frequency of transitions between states; and FO entropy, a measure of variability in the state visits, where high entropy indicates balanced state visits and zero indicates that only one state is visited. Another possibility is to test specific parameters of the HMM, such as transition probabilities or specific state parameters.
For the purposes of this protocol, any form of data $D$, whether it is continuous, epoched, or aggregated, is considered to originate from the HMM and is referred to as ”brain data.” However, $D$ does not need to be a product of the HMM; it can represent any measure as long as it has the correct structure. When $D$ is structured as a 3D tensor with dimensions [No. of timepoints $\times$ No. of subjects or sessions $\times$ No. of states or features], statistical tests can be performed for each timepoint to analyse the temporal aspects of the data. Alternatively, if $D$ is structured as a 2D matrix with dimensions [No. of subjects or sessions $\times$ No. of channels or features], statistical testing is performed on temporally aggregated data. The behavioural matrix $R$, which will be tested against $D$ as shown in Part 1 of Figure 2, has dimensions [No. of subjects or sessions $\times$ No. of behavioural features]; these features can include any non-imaging variables such as cognitive capacity, age, sex, or the experimental condition.
\@fb@secFB
Statistical Analysis
All the settings needed to perform these statistical tests are listed in Table 1. These, unless obvious, will be explained next.
Types of tests Across-subjects
Across-trials &
Across-sessions
Across-state-visits
Input data
- Brain data (D)
- Behavioural data (R)
- Viterbi path (D)
- Non-imaging signals (R)
Permutation/
Parametric testing
Parametric testing if number of permutations set to 0
Methods supported
- multivariate
- univariate
- CCA
- multivariate
- univariate
- CCA
- one-state-vs-the-rest
- one-state-vs-another-state
Category identification
Automatically detects data type in (R) and columnwise applies:
- Independent t-test for boolean values
- ANOVA (F-test) for categorical variables
- F-regression for continuous variables in multivariate tests
- Pearson’s correlation t-test for continuous variables in univariate tests
Default: False
Test combination
Supports NPC on p-values across rows, columns, or both
Default: False
Confounding variables
Regresses out confounding effects from matrix D & matrix R
Default: None
Handling subject
dependencies
Hierarchical permutations for family relationships N/A
Multiple testing correction & Cluster statistics
Supports corrections such as Bonferroni, Benjamini-Hochberg,
FWER (e.g., MaxT), and cluster-based statistics (for spatial or temporal dependencies).
Output
Returns a dictionary with p-values, base statistics,
test types, methods used, and correction details
Table 1: Options for statistical testing
\@fb@secFB
Statistical Analysis: Step 6 - Types of tests
Permutation testing, the primary method used in the framework, is a non-parametric approach that shuffles the data to generate a surrogate distribution where the key property we want to test (and not others) is broken. However, the toolbox also allows for parametric testing, which is computationally much faster and applicable on small samples and when assumptions hold reasonably.
Another feature of the framework is its ability to handle missing values (NaNs) in the dataset. During the analysis, tests automatically exclude these missing values so that incomplete data does not interfere with calculations or affect result reliability. This approach assumes that missing values occur completely at random (MCAR). If the missingness follows other patterns, such as missing at random (MAR) or missing not at random (MNAR), this approach may produce biased or even invalid results, depending on the extent and nature of the missing data.
As stated, this protocol supports the following types of tests: across-subject, across-trials, across-sessions within-subject, and across-state-visits.
\@fb@secFB
Statistical Analysis: Step 7 - Methods supported
The choice of method depends on the specific research question and data structure. The statistical measures used to assess the relationship between the matrices $D$ and $R$ derive from different functions (e.g. regression metrics and correlation coefficients). In all cases, the null hypothesis is that there is no association between the brain data $D$ and the behavioural data $R$. Within this framework, $D$ is structured as $N\times p$, and $R$ is structured as $N\times q$, where:
  • $N$: Number of observations (e.g., subjects or trials),
  • $p$: Number of predictors (e.g., features in $D$),
  • $q$: Number of outcomes (e.g., behavioural variables in $R$) being tested.
For across-subjects, across-session-within-subject, and across-trials-within-session tests, the protocol provides multivariate regression tests and univariate tests as well as Canonical Correlation Analysis (CCA). See Part 2 in Figure 2.
Multivariate regression tests examine the overall relationship between $D$ and each variable or outcome in $R$. This approach produces p-values, one for each outcome in $R$. For example, if $R$ represents 12 HMM states and $R$ includes two behavioural variables like sex and age, the output contains two p-values, one for each behavioural variable. The setup can also be reversed by treating $R$ as the independent variable and $D$ as the dependent variable, with the former case being the default. Multivariate tests use F-statistic as the default base statistic for permutation testing. To assess the predictors’ contribution to the prediction, it also returns regression coefficients and individual p-values per regressor or predictor (similar to those derived from t-tests in multiple linear regression).
Univariate tests independently assess the relationship between each feature in $D$ and each variable in $R$. Using the same example, the output is a $12\times 2$ matrix of p-values, where each element reflects the association between a specific predictor in $D$ and an outcome in $R$. The default base statistic for univariate tests is the t-statistic derived from Pearson correlation.
CCA provides a single p-value summarising the overall relationship between the variables in $D$ and $R$, capturing how brain states in $D$ relate to the behavioural measures in $R$. By default, the analysis includes 1 CCA component, but users can specify a different number of components if desired.
For across-state-visits tests, where we assess the relationship between state time courses and another simultaneously collected set of time series, the protocol includes two additional methods: one-state-vs-the-rest (OSR) and one-state-vs-another-state (OSA) contrasts. Here, $D$ is given as the Viterbi path (such that $p$ is the number of states), i.e. the most likely sequence of states over time, with each time point categorically assigned to one state; and $R$ represents e.g. a set of physiological time series (such as pupil size or skin conductance). Assuming that $R$ has a single column for simplicity, in OSR tests the mean value of $R$ for a specific state is compared to the mean value of $R$ across all other states. By default, the test evaluates whether the mean $R$ for the specific state is larger than the average of the remaining states. This produces $p$ p-values. In OSA, the mean values of $R$ are compared between all possible pairs of states, generating a $p$-by-$p$ matrix of p-values. Each comparison is based on the difference in the mean value of $R$ between two states.
Statistical Analysis: Test Combination (optional)
For the Across-subjects, Across-session-within-subject, and Across-trials-within-session tests, the protocol includes the Non-Parametric Combination (NPC) algorithm to combine multiple p-values into fewer p-values with increased statistical power [24, 25]. Specifically, instead of getting a p-value for each pair of variables—that is a ($p\times q$) matrix of p-values—the NPC algorithm condenses these into either one p-value per row ($1\times p$ p-values), or one p-value per column ($1\times q$ p-values), or a single p-value for the entire test. In our implementation, we use Fisher’s method as the combining function, which efficiently aggregates the p-values while maintaining sensitivity to small values. This approach differs from CCA, which also produces a single p-value but does so by testing the strength of a multivariate relationship between two variable sets. In addition to statistical inference, CCA provides a latent representation of the data in the form of canonical variables—linear combinations that maximise correlation between the sets. NPC, by contrast, aggregates test results across multiple comparisons. While it offers flexibility and interpretability, it does not yield a latent representation or model the joint multivariate structure directly.
\@fb@secFB
Statistical Analysis: Step 8 - Multiple testing correction and cluster statistics
When performing statistical tests, we need to correct for multiple testing correction to control false positives or type 1 errors. The protocol includes standard correction methods from the statsmodels module, including Bonferroni and False Discovery Rate (FDR) control using the Benjamini-Hochberg procedure. Additionally, it supports family-wise error rate (FWER) correction with the MaxT method[26].
For data with spatial or temporal structure, the protocol also includes support for cluster-level inference[27]. A cluster is a contiguous group of tests that survive a predefined statistical threshold. Clusters can be formed by multiple neighbouring voxels or consecutive time points, depending on the type of analysis. The test statistic for a cluster can be its size (called cluster extent) or the sum of the test statistics within it (called cluster mass). Significance of each cluster is assessed by comparing its test statistic to the distribution of the maximum test statistic across all clusters. This distribution is obtained through permutation testing, and since it is based on the maximum statistic, the p-values are FWER-corrected for multiple testing at the cluster level.
\@fb@secFB
Visualisation: Step 9 - Visualising statistical results
Finally, the protocol includes steps for visualising and interpreting results in a way that is both clear and easy to understand, using various graphical tools like heatmaps, bar graphs, and line plots to display the p-values; Part 3 in Figure 2 shows some examples. To highlight significant differences, we use a colour map in a logarithmic scale that shifts from dark red to yellow where there is significance, and from grey to blue where there is not.
Comparison with other methods
This work presents a host of tools for statistical testing of brain dynamics in neuroimaging and neurophysiology. Unlike mainstream toolboxes like FSL, SPM, FreeSurfer, and MNE, whose functionality for statistical testing primarily focus on applying General Linear Models (GLMs) for traditional task-based designs, this protocol supports a wider range of experimental designs, including resting-state paradigms. While solutions for testing in resting-state paradigms exist in other toolboxes, such as functional connectivity approaches (FSL, SPM), surface-based connectivity (FreeSurfer), and spectral analyses (MNE), these approaches typically rely on time-averaged statistics (“static” models) and do not account for temporal variability. Our protocol directly addresses this limitation by enabling dynamic analyses that capture time-varying brain activity and its relationship with behavioural or physiological measures.
Across-subjects and Across-trials analyses are available in FSL, SPM, FreeSurfer, and MNE to perform group-level comparisons and the testing of experimental conditions. However, they do not support across-sessions analyses for studies tracking changes over time, such as longitudinal studies. This protocol uniquely supports Across-sessions testing, as well as across-state-visits testing. The latter examines relationships between dynamic brain states and concurrently measured signals, such as physiological data, using methods like OSR and OSA.
All of the aforementioned toolboxes support both univariate and multivariate tests. However, while FSL, SPM, and FreeSurfer can extend their functionality to include CCA through external tools like Permutation Analysis of Linear Models (PALM), this requires additional setup. MNE lacks this capability entirely. In contrast, CCA is fully integrated into this toolbox, making it easier to use without additional setup.
Our framework also provides hierarchical permutation testing to account for family relationships between subjects. This functionality is fully integrated into the code using PALM, with no requirement for external input or additional installation. MNE does not directly support hierarchical testing.
Regarding multiple testing correction and cluster statistics, all tools support standard approaches, including FWER and FDR. However, this protocol incorporates these corrections directly into its workflow, to streamline their application. Table 2 summarises the main statistical testing capabilities of this protocol in comparison with FSL, SPM, FreeSurfer, and MNE.
Category Feature FSL SPM FreeSurfer MNE-Python
Type of Tests across-subjects Yes Yes Yes Yes
across-trials Yes Yes Yes Yes
across-sessions
within-subject
No No No No
across-state-visits No No No No
Methods Supported Multivariate Yes (via randomize tool) Yes (e.g., MANOVA) Yes Yes
Univariate Yes Yes Yes Yes
CCA
No (default);
Yes via PALM
No (default);
Yes via PALM
No (default);
Yes via PALM
No
one-state-vs-the-rest No No No No
one-state-vs-another-state No No No No
Hierarchical Permutation
Account for family
structure
No (default);
Yes via PALM
No (default);
Yes via PALM
No (default);
Yes via PALM
No
Multiple testing correction & Cluster statistics
Family-Wise Error Rate
Yes Yes Yes Yes
False Discovery Rate
Yes Yes Yes Yes
Cluster-based statistics Yes Yes Yes Yes
Table 2: Comparison of the statistical testing capabilities of GLHMM with FSL, SPM, FreeSurfer, and MNE. Each row represents a specific feature or analysis type, with ’Yes’ indicating that the toolbox supports the functionality and ’No’ indicating that it does not support the functionality.
Expertise needed to implement the protocol
The toolbox described in this protocol is designed to be easy to use for practitioners with varying levels of programming experience, although some basic familiarity with Python is required. While extensive expertise in statistical methods is not required, users should have some ability to interpret the results appropriately. To facilitate its application, the protocol includes clear documentation, example datasets, and tutorials for each of the four statistical test designs. Each tutorial has step-by-step instructions with practical examples, so users do not have to write code from scratch. This allows users to learn how to train an HMM model, select the appropriate data for input, interpret the results, and draw meaningful conclusions with minimal time investment.
Limitations
The presented toolbox exclusively uses linear models for statistical testing. This can be seen as a limitation when the relationships in the data are nonlinear. However, the linear methods presented can easily be extended to the nonlinear case, by using an appropriate basis expansion[28].
Materials
Data
This protocol outlines pipelines for a comprehensive set of statistical tests, applicable to a broad range of scientific questions in neuroscience. We demonstrate these tests using publicly available data, as summarised below. In Protocol 1 (across-subjects), we analyse resting-state brain activity from 1,001 Human Connectome Project (HCP) participants across four sessions to examine its relationship to 15 traits related to cognitive performance, refer to Table S1. In Procedure 2 (across-trials), we study MEG data from a single person who participated in 15 sessions. During each session, the person watched both animate and inanimate objects while their brain activity in the occipital lobe was recorded. This analysis assesses differences in the brain responses when the person looks at animate objects compared to inanimate ones. In Procedure 3 (across-sessions within-subject), we use the same dataset as in Procedure 2, but this time focus on changes over multiple sessions. This analysis shows whether the person exhibits changes in stimulus processing over time (i.e. across sessions) due for example, to learning, or if their brain representations remain stable. In Procedure 4 (across-state-visits), we analyse MEG data from 10 subjects scanned at rest in a dark room. During the scans, pupil size and brain activity were measured concurrently. Nine subjects completed two sessions, and one completed a single session. This analysis explores how changes in brain states, like the default mode network, relate to variations in pupil size[29, 30]. All data (except HCP) needed to reproduce the results of these workflows are hosted on Zenodo (DOI: 10.5281/zenodo.15213970), while the code is available at GitHub.
Software
  • Computer requirements: Any PC, Mac, or Linux computer can be used to run this protocol. $\blacktriangle$ CRITICAL If a local computer is unavailable, the protocol can run via Google Colab for free. For this, a computer with a stable internet connection and a modern web browser such as Chrome or Firefox, is required.
  • Python installation: Download and install Python from the official website: https://www.python.org/downloads/. $\blacktriangle$ CRITICAL Make sure to install the version compatible with the GLHMM package requirements. See https://github.com/vidaurre/glhmm.
  • Recommended tools: To manage Python packages and environments effectively, we recommend using Anaconda, Spyder or Visual Studio Code. Anaconda simplifies package management and environment setup, while Visual Studio Code provides a robust development environment with useful extensions.
  • GLHMM Python package: Install the GLHMM Python package and its dependencies using pip. The package is available for download at https://github.com/vidaurre/glhmm. This protocol is based on the latest release of GLHMM (version 1.0.0, released on April 7, 2025).
Input Data
  • Temporal brain data: This continuous data can correspond to any neuroimaging modalities, such as fMRI, MEG, EEG, ECoG, or LFP. $\blacktriangle$ CRITICAL Other types of temporal (or, more generally, sequential) datasets can also be used as previously mentioned. Except for the initial steps related to the fitting of the HMM, the across-subject tests can also be applied to structural (i.e. non-functional) data.
  • Behavioural measures: These are additional inputs used in statistical analysis alongside the temporal data. These behavioural measures, or more generally non-imaging traits, can include both discrete and continuous variables, such as age, sex, gender, race, ethnicity, IQ, demographic information, and experimental events like responses to stimuli relevant to the study.
Procedure
$\blacktriangle$
CRITICAL This section provides step-by-step instructions for setting up the Python environment and downloading the required datasets for running the four protocols outlined in this protocol.
  1. Install and set up Python environment Timing: 1–5 min
    To set up Python and manage the required tools and packages, we recommend creating a conda environment. If Anaconda is not installed, it can be downloaded from the Anaconda website. After installing Anaconda, create and activate a new environment by running the following commands:
    conda create --name glhmm_env python=3.9
    conda activate glhmm_env
    This creates an environment with Python version 3.9 or newer. Once the environment is activated, clone the GitHub repository and install the GLHMM package using:
    pip install git+https://github.com/vidaurre/glhmm
    The GitHub repository includes all the code required to run the GLHMM package, along with helpful examples and resources. It also contains a requirements.txt file listing additional tools and packages. Run the following command to install these:
    pip install -r requirements.txt
    Datasets and Jupyter notebook tutorials for each of the four procedures outlined in this paper can be downloaded from Zenodo. The data is available under the following DOI: ’10.5281/zenodo.15213970’. You can manually download the data from this DOI and save it in the same folder as the procedures. Alternatively, the provided Jupyter notebooks are designed to automatically download the data if it is not already present in the expected directories.
    $\blacktriangle$
    CRITICAL After setting up the Python environment, load the necessary libraries required to run the protocols:
    import numpy as np
    import matplotlib.pyplot as plt
    import pickle
    from pathlib import Path
    from glhmm import glhmm, graphics, statistics, io, preproc
    These libraries provide tools for data loading, preprocessing, statistical analysis, and visualisation.
\@fb@secFB
Procedure 1: Across-subject testing
  1. 1.
    Load and prepare data: Load data into Python environment Timing: 0–2 min For Procedure 1, we use data from the HCP Young Adult study[31]. Specifically, we work with resting-state fMRI (rfMRI) data from 1,001 subjects, each with 50 parcellations/channels identified through Independent Component Analysis (ICA).
    Each subject’s data contains 4,800 timepoints across four sessions (1,200 timepoints per session, each lasting about 15 minutes). This data is stored in data_measurement_HCP.npy and loaded into the variable D_raw, referred to as matrix $D$.
    Behavioural data for 15 cognitive traits related to fluid intelligence is stored in data_cognitive_traits_HCP.npy and loaded into R_data, referred to as matrix $R$. Look at Table S1 for the full list of traits. Additional confounds, such as sex, age, and fMRI head motion, are included in confounds_HCP.npy.
    We load the data using Numpy:
    D_raw = np.load("data_HCP/data_measurement_HCP.npy")
    R_data = np.load("data_HCP/data_cognitive_traits_HCP.npy")
    confounds = np.load("data_HCP/data_confounds.npy")
    In summary:
    • $D$: [4800, 1001, 50] - 4800 timepoints, 1001 subjects or sessions, and 50 features..
    • $R$: [1001, 15] - 1001 subjects and 15 features (e.g., cognitive traits).
    • Confounds: [1001, 3] - 1001 subjects and 3 confounding variables.
    $\blacktriangle$
    CRITICAL: If using your own data, make sure that it is structured similarly and stored as Numpy arrays.
  2. 2.
    Load and prepare data: Data structuring for the HMM Timing: $<10\text{s}$.
    When training the HMM, we need to format the data as a 2D matrix of shape [(No. of timepoints $\boldsymbol{\cdot}$ No. of subjects) $\times$ No. of features]. This means combining the timepoints from all subjects into one continuous sequence while keeping features (e.g., brain parcellations in this case) in the second dimension.
    For example, our dataset $D$ is shaped like [4800, 1001, 50] ([No. of timepoints $\times$ No. of subjects $\times$ No. of features]), it needs to be reshaped to [4804800, 50].
    We use the get_concatenate_subjects function to reshape the data by concatenating the timepoints, and the get_indices_timestamp function to create indices marking where each subject’s data starts and ends:
    D_con = statistics.get_concatenate_subjects(D_raw)
    idx_subject = statistics.get_indices_timestamp(
    D_raw.shape[0],
    D_raw.shape[1])
    The generated indices for each session will look like this:
    [[ 0 4800]
    [ 4800 9600]
    [4800000 4804800]]
  3. 3.
    Load and prepare data: Preprocessing data Timing: $<10\text{s}$
    Since the data from the HCP is already cleaned and preprocessed when downloaded, the next step is to standardise the full time series before performing further analysis.Standardising is important, especially when comparing data between different individuals as it helps ensure that the analysis is not affected by noise or differences in measurement scales.
    To standardise the data, we use the preprocess_data function from the module preproc. This function has many options for processing data, but here we focus on standardisation. Standardisation makes sure each signal has an average (mean) of zero and a standard deviation of one.
    Here is how to do it:
    D_preproc, idx_preproc = preproc.preprocess_data(
    D_con,
    idx_subject,
    standardise=True
    By entering the concatenated data (D_con) and the indices of each subject (idx_subject) and setting standardise=True, the function standardises the whole dataset.
  4. 4.
    Load and prepare data: Set-up and train a HMM Timing: $<4\text{hours}$.
    To start using GLHMM, the first step is to set up the GLHMM model and choose the right settings. For a standard Gaussian HMM, we are not focused on interactions between different sets of data, so we set model_beta=’no’.
    In this example, the number of states is controlled by the parameter K and is set to 12. Each state is represented as a Gaussian distribution with its own unique average (mean) and full covariance structure, meaning each state has a distinct pattern. To set this up, we set covtype=’full’, and the model handles the state-specific mean automatically.
    Here is how to initialise the GLHMM model:
    K = 12
    hmm_HCP = glhmm.glhmm(model_beta=’no’, K=K, covtype=’full’)
    Once the model is initialised, it is time to train the HMM using the preprocessed data D_preproc and the subject index matrix idx_subject. In this case, we are not modelling interactions between two different time series, so we set X=None. The Y input should be the preprocessed time series data (D_preproc) that we want to use for estimating states.
    Gamma,Xi,FE = hmm_HCP.train(X=None, Y=D_preproc, indices=idx_subject)
    The trained model returns Gamma (the state probabilities at each timepoint), Xi (the joint probabilities of past and future states conditioned on the data), and FE (the free energy of each iteration).
  5. 5.
    Load and prepare data: Configure HMM outputs for statistical analysis Timing: $<40\text{s}$.
    To prepare for statistical analysis, we calculate the aggregated summary statistics from the Gamma values. For each subject, we compute FO, which represents the probability distribution of time spent in each state. FO shows how much time a subject spends in each of the 12 states across their recording. These values are our brain data input for the statistical test for each subject and are stored in the variable D_fo, which is generated using the following code:
    D_fo = glhmm.utils.get_FO(Gamma, idx_subject)
    The resulting matrix has dimensions [1001, 12], where each row corresponds to a subject and each column represents a state. The values in each row sum to one and provide a normalised summary of the time spent in each state.
  6. 6.
    Statistical Analysis: Types of test Timing: $<1\text{s}$.
    For statistical testing, we use the test_across_subjects function from the statistics.py module to test the relationship between D_fo (brain data, $D$) and R_data (behavioural measurements, $R$) for each subject. $\blacktriangle$ CRITICAL The test_across_subjects function assumes that all subjects can be permuted without affecting the results, known as being exchangeable. However, in practice, some subjects may be related, which violates this assumption. To handle this, we need to use an Exchangeability Block (EB) file to organise subjects into family blocks such that any permutations of the data respect family structures.
    Creating the EB.csv File The EB.csv file organises subjects into family blocks to maintain these structures during permutation testing. You can specify the file location as shown below:
    # Exchangeability Block (EB) information
    dict_fam = {
    ’file_location’: ’EB.csv’
    A step-by-step tutorial for creating an EB.csv file for the HCP dataset is available in this notebook[21].
  7. 7.
    Statistical Analysis: Methods supported Timing: $<10\text{min}$
    With the EB.csv ready, the next step is to set-up and run the test_across_subjects function.
    Below is an example of how to configure the required inputs and perform a multivariate statistical test with 10,000 permutations. For this analysis, we use brain data (D_fo) and behavioural measurements (R_data) as inputs. Additional parameters include confounds (confounds_data), the number of permutations (Nnull_samples), the analysis method (method), and the family dictionary (dict_family).
    # Set parameters for multivariate testing
    method = "multivariate"
    Nnull_samples = 10_000 # Number of permutations
    # Perform multivariate analysis
    result_multivariate = statistics.test_across_subjects(
    D_data=D_fo,
    R_data=R_data,
    confounds=confounds_data,
    Nnull_samples=Nnull_samples,
    method=method,
    dict_family=dict_fam,
    The results of the test are stored in the result_multivariate dictionary, which contains detailed results, including p-values, test statistics, and baseline measures. For a breakdown of the dictionary structure, see Box 1.
    Box 1 — Details of the Results Dictionary The result_test dictionary stores the outputs of statistical tests and contains: ’pval’ — P-values computed under the null hypothesis. ’base_statistics’ — The observed test statistic calculated from the original (unpermuted) data. ’null_stat_distribution’ — Test statistics generated under the null hypothesis, where the first row corresponds to the observed test statistic (’base_statistics’). ’statistical_measures’ — Dictionary specifying the type of test statistic in each column in (’base_statistics’), such as t-statistics or F-statistics. ’test_type’ — Type of test performed (e.g., across_subject_test). ’method’ — Analytical approach used (e.g., multivariate and univariate). ’combine_tests’ — Indicates whether the Non-Parametric Combination (NPC) method was applied to summarise p-values. ’max_correction’ — Whether Max-statistic correction was used for multiple comparisons. ’Nnull_samples’ — Total number of null samples including the observed one. ’test_summary’ — Dictionary summarising test results. ’pval_f_multivariate’ — F-test p-values for multivariate tests and Nnull_samples > 1. ’pval_t_multivariate’ — t-test p-values for multivariate tests and Nnull_samples > 1.
  8. 8.
    Statistical Analysis: Multiple testing correction and cluster statistics Timing: $<5\text{s}$
    To minimise the risk of false positives in statistical tests, we can apply FWER correction using the MaxT method. This is done during permutation testing by setting FWER_correction=True in the test_across_subjects function. This adjustment accounts for multiple testing correction by modifying the significance levels. If a different correction method, such as Bonferroni or Benjamini-Hochberg, is preferred, we can use the pval_correction function to adjust the p-values after running the test. In this example, we apply the Benjamini-Hochberg method by providing the p-values from result_multivariate and setting method=’fdr_bh’:
    pval_corrected, rejected_corrected =
    statistics.pval_correction(result_multivariate, method=’fdr_bh’)
    The function returns two outputs: pval_corrected, which contains the adjusted p-values, and rejected_corrected, a boolean array indicating which hypotheses are rejected based on the corrected p-values.
  9. 9.
    Visualisation: Visualising statistical results Timing: $<5\text{s}$
    Visualising both uncorrected and corrected p-values helps identify significant findings before and after applying corrections. Figure 3 displays results from multivariate and univariate tests. Although this demonstration focuses on multivariate testing, running a univariate test is straightforward. Set method="univariate" when running the test. The visualisations include bar plots for multivariate tests and heatmaps for univariate tests. These functions are part of the graphics module. For the different plots, we use alpha=0.05 to highlight p-values below this threshold:
    # Features of cognitive traits - see Table S1 for details.
    features = [’ReadEng_Unadj’,’PicVocab_Unadj’,’WM_Task_Acc’,
    ’PMAT24_A_CR’,’ListSort_Unadj’,’PMAT24_A_SI’,’PicSeq_Unadj’,
    ’VSPLOT_TC’,’CardSort_Unadj’,’Language_Task_Acc’,’Flanker_Unadj’,’IWRD_TOT’,’ProcSpeed_Unadj’,’PMAT24_A_RTCR’, ’MMSE_Score’
    alpha = 0.05 # Threshold for the p-value plots
    graphics.plot_p_values_bar(
    result_multivariate["pval"],
    title_text="Multivariate Test - Uncorrected",
    alpha=alpha,
    xticklabels=features,
    xlabel_rotation=45,
    # Plot corrected p-values
    graphics.plot_p_values_bar(
    pval_corrected,
    title_text="Multivariate Test - Benjamini-Hochberg",
    alpha=alpha,
    xticklabels=features,
    xlabel_rotation=45,
    The bar plots display p-values for each feature, with corrected values shown using the Benjamini-Hochberg method. For univariate tests, use the plot_p_value_matrix function to generate a heatmap of p-values.
    Figure 3: Result from procedure 1. a) Results of the multivariate tests without and with the Benjamini-Hochberg method. b) Results of the univariate tests with the same settings.
\@fb@secFB
Procedure 2: Across-trials testing
  1. 1.
    Load and prepare data: Load data into Python environment Timing: $<1\text{min}$
    $\blacktriangle$
    CRITICAL For Procedure 2, we analyse MEG data collected from a single person who participated in 15 experimental sessions over approximately six months. During each session, the participant engaged in multiple trials, and viewed animate and inanimate objects.
    The data includes 72 MEG channels around the occipital lobe, stored in data_raw.pkl. This is a list where each element corresponds to a session. These data is loaded into the variable D_raw, referred to as matrix $D$.
    Behavioural data for the 15 sessions is stored in R_data.pkl and loaded into the variable R_data, referred to as matrix $R$. This list contains trial-level information indicating whether the stimulus presented was an animate or inanimate object. This enables testing whether the brain responds differently to these conditions.
    To perform epoch-based analyses, we also load a list of event markers stored in event_markers.pkl. These markers include information about the timepoints at which stimuli were presented and are loaded into the variable event_markers.
    We use the pickle module to load the data:
    with open("data_raw.pkl", "rb") as f:
    D_raw = pickle.load(f)
    with open("R_data.pkl", "rb") as f:
    R_data = pickle.load(f)
    with open("event_markers.pkl", "rb") as f:
    event_markers = pickle.load(f)
    In summary:
    • $D$: MEG data as a list of 2D matrices, where each matrix represents a session and has the shape [No. of timepoints $\times$ No. of channels] for that session.
    • $R$: Behavioural data as a list of arrays, where each array corresponds to a session and encodes trial information (0 for inanimate objects and 1 for animate objects).
    • Event markers: A list of arrays, where each array corresponds to a session and includes the stimulus presentation timepoints and metadata.
  2. 2.
    Load and prepare data: Data structuring for the HMM Timing: $<10\text{s}$.
    For HMM training, we must format the MEG data to a 2D matrix with the shape [(No. of timepoints across all sessions) $\times$ No. of channels]. This involves concatenating the data from all 15 sessions into a single matrix, where each row corresponds to a timepoint and each column represents a MEG channel.
    The function get_indices_from_list is used to generate indices marking where each session starts and ends within the concatenated data.
    Run the following commands to structure the data:
    D_con = np.concatenate(D_raw, axis=0)
    idx_subject = statistics.get_indices_from_list(D_raw)
    The generated indices for each session will look like this:
    [[ 0 1530001]
    [ 1530001 3034002]
    [10496514 10864515]]
  3. 3.
    Load and prepare data: Preprocessing data Timing: $<6\text{mins}$
    In this example, we isolate brain activity in the alpha band (8–13 Hz) to focus on specific oscillatory patterns associated with attention and sensory suppression [32, 33]. Preprocessing involves several steps:
    Preprocessing Steps:
    • Band-pass Filtering: Apply a band-pass filter to extract the alpha frequency band (8–13 Hz). The same procedure can be applied to other frequency bands.
    • Standardisation: Normalise the data to zero mean.
    • Hilbert Transform: Use the Hilbert transform to extract the amplitude (strength) and phase (timing) of brain waves.
[…]
    • Downsampling: Reduce the sampling rate from 1000 Hz to 250 Hz to decrease computational load.
    Run the following code to preprocess the concatenated MEG data (D_con) and corresponding indices (idx_data).
    # Define preprocessing parameters
    freqs = (8, 13) # Alpha band
    pca_variance = 0.9 # Retain 90% variance
    fs = 1000 # Original sampling rate
    f_target = 250 # Target sampling rate after downsampling
    standardise = True # Standardise the data
    onpower = True # Hilbert transform
    # Preprocess the data
    D_preproc, idx_preproc = preproc.preprocess_data(
    data=D_con,
    indices=idx_data,
    fs=fs,
    standardise=standardise,
    filter=freqs,
    onpower=onpower,
    pca=pca_variance,
    downsample=f_target
    After preprocessing, the data is stored in D_preproc, and the corresponding indices are stored in idx_preproc.
  4. 4.
    Load and prepare data: Set-up and train a HMM Timing: $<45\text{mins}$.
    We use a standard Gaussian HMM to identify distinct brain states and track changes over time. The key output, Gamma, provides the probability of being in each state at every timepoint and forms the basis for subsequent analyses.
    The HMM is set up with the same parameters as in Procedure 1, but with 5 states:
    K = 5
    hmm_classic = glhmm.glhmm(model_beta=’no’, K=K, covtype=’full’)
    Train the HMM using the preprocessed MEG data (data_session_preproc) and session indices (idx_data_preproc):
    Gamma, _, _ = hmm_classic.train(
    X=None,
    Y=data_session_preproc,
    indices=idx_data_preproc
    The Gamma matrix contains the probability of each state at every timepoint.
  5. 5.
    Load and prepare data: Configure HMM outputs for statistical analysis Timing: $<40\text{s}$.
    With the Gamma values, we can analyse how brain states relate to specific events in the data. The Gamma matrix has the following dimensions:
    • Gamma: [2716140, 5] - Downsampled from 10,864,515 to 2,716,140 timepoints, with 5 brain states.
    The reduction in timepoints reflects the downsampling from 1000 Hz to 250 Hz..
    Epoch the data To analyse responses to specific events (e.g., stimulus presentations), the Gamma data is divided into smaller segments called epochs. Each epoch corresponds to a trial, defined using event markers. The event marker timestamps need to be downsampled to match the Gamma data, by setting fs_target to 250 Hz. The window length for each epoch is set to 250 timepoints, representing a 1-second time window following the stimulus.
    Execute the following commands to extract the epochs:
    fs_target= 250 # Define the target sampling frequency
    epoch_window_tp = 250 # Epoch window length in timepoints
    # Extract epochs for the HMM state time courses
    gamma_epoch, idx_data_epoch, R_data_epoch = statistics.get_event_epochs(
    Gamma, # State probabilities
    idx_data, # Session indices
    R_data, # Behavioural data
    event_markers, # Event markers
    fs=fs, # Sampling freq
    fs_target=fs_target, # Downsampled freq
    epoch_window_tp=epoch_window_tp # Epoch window
    The resulting dimensions are:
    • gamma_epoch: [250, 8368, 5] — 250 timepoints per trial, 8368 trials, and 5 states.
    • R_data_epoch: [8368] — Stimulus labels for each trial (0 for inanimate, 1 for animate).
    • idx_data_epoch: Marks the start and end trial indices for each session.
    This configuration prepares the Gamma data for statistical analyses by segmenting it into epochs aligned with the experimental events.
  6. 6.
    Statistical Analysis: Types of test Timing: $<1\text{s}$.
    We use the test_across_trials function from the statistics.py module to test whether the brain states (gamma_epoch, $D$) process the behavioural conditions (R_data_epoch, $R$) — watching animate vs. inanimate objects — in the same way for each trial, or whether the responses vary across trials.
  7. 7.
    Statistical Analysis: Methods supported Timing: $<20\text{min}$
    Below, we show how to configure the required inputs and perform a multivariate statistical test with 10,000 permutations.
    For this analysis, we use brain data (gamma_epoch) and behavioural conditions (R_data_epoch) as inputs. Additional parameters include the indices for each session (idx_data_epoch), the number of permutations (Nnull_samples), and the analysis method (method)
    Run the following code:
    # Set parameters for multivariate testing
    method = "multivariate"
    Nnull_samples = 10_000 # Number of permutations
    # Perform across-trial testing
    results_multivariate = statistics.test_across_trials(
    D_data=gamma_epoch,
    R_data=R_data_epoch,
    idx_data=idx_data_epoch,
    Nnull_samples=Nnull_samples,
    method=method
    The results of the test are stored in the result_multivariate dictionary. For a breakdown of the dictionary structure, see Box 1 in Procedure 1. $\blacktriangle$ CRITICAL The option test_statistics_option=True is required only if cluster-level inference will be used during multiple testing correction at Step 8. By default, it is set to True.
  8. 8.
    Statistical Analysis: Multiple testing correction and cluster statistics Timing: $<5\text{s}$
    In this example, we demonstrate how to perform cluster-level inference, which identifies clusters of significant results while reducing the risk of false positives. The correction uses the output from result_multivariate, and the test focuses on p-values below a threshold of 0.01, specified by alpha=0.01. The pval_cluster_based_correction function performs the correction:
    pval_cluster = statistics.pval_cluster_based_correction(
    results_multivariate,
    alpha=alpha
    The function returns pval_cluster, which contains the adjusted p-values after cluster-level inference. $\blacktriangle$ CRITICAL Alternative multiple testing correction procedures, such as the MaxT method (FWER correction), Bonferroni, or Benjamini-Hochberg can also be used.
  9. 9.
    Visualisation: Visualising statistical results Timing: $<5\text{s}$
    This step visualises the results for both uncorrected and cluster-corrected p-values from the multivariate test, as shown in Figure 4. The function plot_p_values_over_time from the graphics module is used to generate line plots. For this example, alpha=0.01 is set to highlight p-values below this threshold.
    # Set parameters
    xlabel = "Time (ms)"
    alpha = 0.01
    # Plot uncorrected p-values
    graphics.plot_p_values_over_time(
    results_multivariate["pval"],
    title_text=f"Uncorrected - Alpha",
    xlabel=xlabel,
    alpha = alpha,
    # Plot cluster corrected p-values
    graphics.plot_p_values_over_time(
    pval_cluster,
    title_text=f"Cluster Corrected - Alpha",
    xlabel=xlabel,
    alpha=alpha,
    Figure 4: Result from Procedure 2. a) Average state probabilities over time for inanimate and animate stimuli, and the difference between the two conditions. Significant differences after cluster-level inference are highlighted in grey. b) Results of the multivariate test (uncorrected). c) Results of the multivariate test after applying cluster-level inference.
\@fb@secFB
Procedure 3: Across-session-within-subject testing
Steps 1–5 follow the same process as in Procedure 2, as this protocol uses the same data and preprocessing setup. For details on data loading, preparation, and HMM training, refer to Procedure 2. This protocol starts at Step 6 and focuses on the statistical analysis for the across-sessions within-subject test.
  1. 6.
    Statistical Analysis: Types of test Timing: $<1\text{s}$.
    We use the test_across_sessions_within_subjects function from the statistics.py module to test whether the brain states (gamma_epoch, $D$) encode the behavioural conditions (R_data_epoch, $R$) — watching animate vs. inanimate objects — consistently across sessions or if the encoding changes. Differences may suggest that the brain processes the same task differently across different sessions over time.
  2. 7.
    Statistical Analysis: Methods supported Timing: $<17\text{min}$
    Below, we show how to configure the required inputs and perform a multivariate statistical test with 10,000 permutations. For this analysis, we use brain data (gamma_epoch) and behavioural conditions (R_data) as inputs. Additional parameters include the indices for each session (idx_data_epoch), the number of permutations (Nnull_samples), and the analysis method (method).
    Run the following code:
    # Set parameters for multivariate testing
    method = "multivariate"
    Nnull_samples = 10_000 # Number of permutations
    # Perform across-trial testing
    results_multivariate = statistics.test_across_sessions_within_subject(
    D_data=gamma_epoch,
    R_data=R_data_epoch,
    idx_data=idx_data_epoch,
    Nnull_samples=Nnull_samples,
    method=method
    The results of the test are stored in the result_multivariate dictionary. For a breakdown of the dictionary structure, see Box 1 in Procedure 1.
  3. 8.
    Statistical Analysis: Multiple testing correction and cluster statistics Timing: $<5\text{s}$
    In this example, we demonstrate how to apply multiple testing correction by using FWER correction with the MaxT method. To run the test, we use the function pval_FWER_correction, and it requires only the result_multivariate dictionary as input.
    pval_FWER =
    statistics.pval_FWER_correction(result_multivariate)
    The function returns the FWER-corrected p-values in the variable pval_FWER.
  4. 9.
    Visualisation: Visualising statistical results Timing: $<5\text{s}$
    The results for uncorrected, FWER-corrected, Benjamini-Hochberg, and cluster-corrected p-values are shown in Figure 5 for both multivariate and univariate tests. While the code example and text focus on the multivariate test with FWER correction, the figure provides a broader overview of different correction methods.
    Notably, since this dataset includes only one variable (stimulus presentation of animate and inanimate objects), FWER correction has no effect on the multivariate test results. This is expected, as MaxT correction only applies when multiple tests are performed. However, in the univariate test, FWER correction does show an effect, as the permutation process involves multiple tests across time points.
    To perform a univariate test, set method="univariate" in the statistical testing function. Line plots visualise multivariate tests, and heatmaps are used for univariate tests. Both methods are part of the graphics module.
    For this example, alpha=0.01 is set to highlight p-values below the threshold.
    # Set up parameters
    xlabel = "Time (ms)"
    # Threshold for the p-value plots
    alpha = 0.01
    # Plot uncorrected p-values
    graphics.plot_p_values_over_time(
    results_multivariate["pval"],
    title_text=f"Uncorrected - Alpha",
    xlabel=xlabel,
    alpha = alpha,
    # Plot FWER corrected p-values
    graphics.plot_p_values_over_time(
    pval_FWER,
    title_text=f"FWER - Alpha",
    xlabel=xlabel,
    alpha = alpha
    Figure 5: Result from procedure 3. a) Multivariate test results for uncorrected, FWER-corrected, Benjamini-Hochberg, and cluster-corrected p-values. b) Univariate test results for the same correction methods.
\@fb@secFB
Procedure 4: Across-state-visits testing
  1. 1.
    Load and prepare data: Load data into Python environment Timing: $<5\text{s}$ For Procedure 4, we analyse data collected from 10 participants during resting-state MEG recordings. Each participant completed two sessions, except for one participant who completed only one session. The MEG data, stored in data_meg.pkl, is a list where each element corresponds to a session. It is loaded into the variable data_meg, referred to as matrix $D$.
    Pupillometry, recorded simultaneously for each session, is stored in pupillometry.pkl and loaded into the variable data_pupillometry, referred to as matrix $R$.
    Unlike in previous protocols, where the HMM was trained on the dataset, we use a pre-trained Temporal Delayed Embedding HMM (TDE-HMM) from Vidaurre et al. 2018[15]. We will use this pre-trained model to decode brain states for each session of MEG data without requiring additional training. The pre-trained model is stored in the MATLAB file hmm.mat. To decode the MEG data:
    • Load the pre-trained TDE-HMM using the read_flattened_hmm_mat function from the io module.
    • Retrieve the model settings using scipy.io.loadmat.
    We use the pickle module to load the MEG and pupillometry data:
    with open("data_meg.pkl", "rb") as f:
    data_meg = pickle.load(f)
    with open("pupillometry.pkl", "rb") as f:
    data_pupillometry = pickle.load(f)
    # Load pre-traind TDE-HMM
    hmm_TDE = io.read_flattened_hmm_mat("hmm.mat")
    # Load the settings of the TDE-HMM
    hmm_TDE_settings = scipy.io.loadmat(’hmm.mat’)
    In summary:
    • data_meg: Brain activity as a list of 19 sessions, where each session is a 2D matrix with shape [No. of timepoints $\times$ No. of channels] (42 channels extracted using PCA).
    • data_pupillometry: Pupil size as a list of 19 sessions, where each session is a 1D array with shape [No. of timepoints].
    • hmm_TDE: Pre-trained TDE-HMM model[15] used to decode brain states from brain data.
    $\blacktriangle$
    CRITICAL Ensure the MEG and pupillometry data are temporally aligned for accurate state decoding and statistical testing. The number of timepoints in data_meg and data_pupillometry must match for each session. Any mismatched data lengths could lead to errors during analysis.
  2. 2.
    Load and prepare data: Data structuring for the HMM Timing: $<5\text{s}$.
    Before applying the pre-trained TDE-HMM model, we need to organise the MEG data to a 2D matrix with the shape [(No. of timepoints across all sessions) $\times$ No. of channels]. This involves concatenating the data from all 19 sessions into a single matrix, where each row corresponds to a timepoint and each column represents a MEG channel.
    The pupil size data (data_pupillometry) also needs to be structured along the time dimension to form a single 1D array: [No. of timepoints across all sessions].
    To track the start and end timepoints for each session, we generate an index matrix using the function get_indices_from_list. The resulting matrix has the shape [No. of sessions $\times$ 2], where each row specifies the start and end timepoints for a session.
    Run the following commands to structure the data:
    D_con = np.concatenate(data_meg, axis=0)
    R_data = np.concatenate(data_pupillometry, axis=0)
    idx_data = statistics.get_indices_from_list(data_meg)
    The generated indices for each session will look like this:
    [[ 0 85996]
    [ 85996 167903]
    [10496514 1473460]]
  3. 3.
    Load and prepare data: Preprocessing data Timing: $<20\text{s}$
    Before analysing the data using TDE-HMM, the MEG data must be formatted correctly. This preparation involves two main steps:
    • Preprocess the brain data (D_preproc)
      We will standardise the data to ensure that all time series data are on the same scale to ensure comparability. This step uses the preprocess_data function to standardise the data to a zero mean and unit variance.
    • Preparing data for the TDE-HMM (D_tde)
      The build_data_tde function prepares the MEG data for TDE-HMM analysis by:
      • Adding time lags to capture changes in brain activity over short windows. For this example, 7 time lags before and after each timepoint are used.
      • Applying a PCA projection to reduce dimensionality, with settings extracted from hmm_TDE_settings.
    We will use the following script to preprocess the data:
    # Preprocess data
    D_preproc, idx_data_preproc = preproc.preprocess_data(
    data=D_con,
    indices=idx_data,
    standardise=True, # Standardise the data
    # Specify time lags
    embedded_lags = 7
    lags = np.arange(-embedded_lags, embedded_lags + 1)
    # Load PCA projection settings
    pca_proj = hmm_TDE_settings["train"]["A"][0][0]
    # Build the MEG data in TDE format
    D_tde, indices_tde = preproc.build_data_tde(
    D_preproc,
    idx_data_preproc,
    lags=lags,
    pca=pca_proj
    Now the MEG data is ready for decoding brain states with the pre-trained TDE-HMM model.
  4. 4.
    Load and prepare data: Set-up and train a HMM Timing: $<1\text{min}$.
    We can use the TDE-HMM to decode brain activity into distinct states over time. These states form a sequence called the Viterbi Path (D_vpath_tde), which shows the brain’s most likely state at each time point. The across-state-visits test is the only statistical test in this framework that relies on the Viterbi Path instead of other outputs, such as state time courses (Gamma).
    We can decode the Viterbi path by using the following command:
    D_vpath_tde = hmm_TDE.decode( X=None, Y=D_tde, indices=indices_tde, viterbi=True )
    $\blacktriangle$
    CRITICAL Ensure that the Viterbi Path (D_vpath_tde) is decoded correctly, as it is the primary input for the across-state-visits test. Errors in decoding or preprocessing may lead to misleading conclusions, so always verify the input data format and preprocessing steps before decoding.
  5. 5.
    Load and prepare data: Configure HMM outputs for statistical analysis Timing: $<40\text{s}$.
    To analyse the relationship between brain states from the Viterbi path (D_vpath_tde, $D$) and pupil size (R_data, $R$), the datasets must be aligned. The dimensions of the data are as follows:
    • Viterbi path (D_vpath_tde): (1473194,12)
    • Pupil size (R_data): (1473460,)
    The difference in length occurs because the TDE-HMM introduces a lag of 7 timepoints at the start and end of each session, and thereby removes 14 timepoints per session.
    To match the dimensions, the Viterbi path is padded to restore the original length of the pupillometry data. This is achieved using the pad_vpath function, which adjusts for the lagged timepoints based on session boundaries.
    embedded_lags = 7
    D_vpath_pad = statistics.pad_vpath(D_vpath_tde, embedded_lags, indices_tde)
    The padded Viterbi path (D_vpath_pad) is initially stored as a 2D array with one-hot encoding, where each row represents a timepoint, and one state is active per row. To simplify the data and reduce memory usage, we convert the array to a 1D format where each value represents the active state for a given timepoint and store the data into the variable D_vpath.
    D_vpath = D_vpath_pad.nonzero()[1] + 1
    $\blacktriangle$
    CRITICAL The alignment of the Viterbi path and pupil size data is essential for performing the statistical testing. Always verify that both datasets match in length after padding. Additionally, plotting FO can provide a useful overview of how consistently the TDE-HMM captures brain activity across sessions as shown in 6a.
  6. 6.
    Statistical Analysis: Types of test Timing: $<1\text{s}$.
    We use the test_across_state_visits function from the statistics.py module to test whether specific brain states (D_vpath, $D$) are associated to differences in pupil size (R_data, $R$) during resting-state recordings.
  7. 7.
    Statistical Analysis: Methods supported Timing: $3-4\text{hours}$
    Across-state-visits analysis includes methods such as OSA and OSR to explore how brain states relate to other signals, like pupil size in our case. Here, we focus on OSA. The OSA test compares pupil size between pairs of brain states. For example, it tests whether the average pupil size during State 1 differs from that in State 2, State 3, and so on. This helps reveal how specific brain states influence pupil size during resting-state recordings.
    To perform this analysis, we use the brain state sequence (D_vpath) and pupil size data (R_data) as inputs. Additional settings include the number of permutations (Nnull_samples) and the test type (method).
    Run the following code to perform the test:
    # Set parameters for the state pair comparison test
    method = "OSA"
    Nnull_samples = 10_000 # Number of Viterbi path surrogates
    # Run the analysis
    results_OSA = statistics.test_across_state_visits(
    D_data=D_vpath,
    R_data=R_data,
    method=method,
    Nnull_samples=Nnull_samples
    The test results are stored in a variable called results_OSA. For a breakdown of the dictionary structure, see Box 2.
    Box 2 — Details of the Results Dictionary ’pval’ — P-values computed under the null hypothesis. ’base_statistics’ — The observed test statistic calculated from the original (unshuffled) Viterbi path. ’null_stat_distribution’ — Test statistics generated under the null hypothesis, where the first row corresponds to the observed test statistic (’base_statistics’). ’statistical_measures’ — Dictionary specifying the type of test statistic in each column in (’base_statistics’), such as t-statistics or F-statistics. ’test_type’ — Type of test performed (across_state_visits). ’method’ — Analytical approach used (e.g., multivariate and univariate). ’max_correction’ — Whether Max-statistic correction was used for multiple comparisons. ’Nnull_samples’ — Total number of Monte Carlo samples (i.e., surrogate Viterbi paths) including the observed one. ’test_summary’ — Dictionary summarising test results. ’pval_f_multivariate’ — F-test p-values for multivariate tests and Nnull_samples > 1. ’pval_t_multivariate’ — t-test p-values for multivariate tests and Nnull_samples > 1.
    $\blacktriangle$
    CRITICAL Creating the permutation matrix is the most time-intensive part of the test because it involves every timepoint in the data. For this example, the matrix has a size of [1473460, 1000]. To save time, you can create this matrix ahead of time and store it (e.g., as vpath_surrogates). Using this precomputed matrix reduces the test runtime to just a couple of minutes.
  8. 8.
    Statistical Analysis: Multiple testing correction and cluster statistics Timing: $<5\text{s}$
    In this example, we apply multiple testing correction using the Benjamini-Hochberg procedure. The function pval_correction performs this correction, and requires the p-values from results_OSA as input, with the method set to ’fdr_bh’.
    # Apply Benjamini-Hochberg correction
    pval_fdr_bh, _ = statistics.pval_correction(
    results_OSA,
    method=’fdr_bh’
    The corrected p-values are stored in the variable pval_fdr_bh.
  9. 9.
    Visualisation: Visualising statistical results Timing: $<5\text{s}$
    For procedure 4, we performed only the OSA test, which compares pupil size between pairs of brain states. However, we also visualise OSR results to provide a reference for both methods. Figure 6 displays the uncorrected and Benjamini-Hochberg corrected p-values for OSA and OSR. For OSA, p-values are stored in a [p, p] array where p = 12(representing 12 states). Values above the diagonal represent comparisons where State $X>$ State $Y$, while values below the diagonal represent State $X<$ State $Y$
    To run an OSR test, set method="OSR" in the statistical testing function. Heatmaps are used for OSA results, while bar plots are employed for OSR. Both visualisation functions are part of the graphics module.
    # Plot uncorrected p-values
    graphics.plot_p_value_matrix(results_OSA["pval"],
    title_text =’OSA - Uncorrected’,
    xlabel="State X",
    ylabel="State Y",
    alpha=0.05,
    none_diagonal=True,
    annot=True,
    x_tick_min=1,
    x_tick_max=12
    # Plot Benjamini-Hochberg corrected p-values
    graphics.plot_p_value_matrix(pval_fdr_bh,
    title_text =’OSA - Benjamini-Hochberg correction’,
    figsize=(9, 5),
    xlabel="State X",
    ylabel="State Y",
    alpha=0.05,
    none_diagonal=True,
    annot=True,
    x_tick_min=1,
    x_tick_max=12
    Figure 6: Results from Procedure 4. a) FO for each session, showing the stability of TDE-HMM decoding. b) OSA test results, with both uncorrected and Benjamini-Hochberg corrected p-values. c) OSR test results, shown for comparison, with the same correction methods applied.
    Visualisation: Spectral and spatial features
    In addition to statistical testing it is useful to explore the spectral and spatial characteristics of the decoded brain states. These features help describe the functional profile of each state and support interpretation of the results. Figure 7 shows various visualisations including power spectra, spectral components from a data driven decomposition (non-negative matrix factorisation or NNMF [15]), and spatial maps of power and coherence of one of the spectral components. All visualisation steps are implemented in the Procedure 4 notebook.
    Figure 7: Spectral and spatial characterisation of brain states. (a) Power spectral density (PSD) profiles for each brain state inferred from the HMM. Highlighted lines indicate states identified as significant by the univariate test. (b) Spectral components obtained by applying non-negative matrix factorisation (NNMF) with four components to the PSDs. (c) Brain map showing the first NNMF component for state 4, showing higher activation in sensory (visual, somatosensory, auditory) areas (d) Connectivity map for state 4 based on the first NNMF component.
Timing
For each step in every procedure, the timing is indicated based on the provided data. These timings were obtained using a modest Lenovo ThinkPad T16 Gen 3 laptop (Intel Core Ultra 7 155U, 32 GB RAM, 1 TB SSD). Of course, the processing time depends on the dataset size, type of test, and applied method, so actual run times may vary.
Anticipated results
The GLHMM framework offers an accessible yet effective set of tools for analysing temporal in any field of research, even though we have focussed here on neuroscience. Using the four presented statistical tests, users can investigate associations between the properties of a dynamic system and a set of external variables. For instance, the Across-subjects test assesses associations between brain states and individual traits or characteristics, while the Across-trials test can pinpoint temporal patterns in experimental conditions. For longitudinal studies benefit, the across-sessions-within-subject test can assesses changes in brain-behaviour relationships over longer time scales. Finally, the across-state-visits test can be used to probe the interactions between brain states and concurrently recorded signals. Each workflow is fully documented in Jupyter notebooks, guiding users through data preparation, analysis, and interpretation, with brief conclusions.
Data availability
All data used in this protocol are freely accessible on Zenodo (DOI: 10.5281/zenodo.15213970). The Zenodo repository includes references to the GitHub repository for associated code.
Code availability
All code is available on GitHub as Python notebooks and can be run for free in the cloud using Google Colab. To ensure reproducibility, the GitHub repository is also archived on Zenodo (DOI: 10.5281/zenodo.15213970). This setup supports versioning to track future updates, including new code and tutorials.
Acknowledgement
We sincerely thank Francesca Fardo for collecting the MEG data used in Procedures 2 and 3. DV is supported by a Novo Nordisk Foundation Emerging Investigator Fellowship (NNF19OC-0054895) and an ERC Starting Grant (ERC-StG-2019-850404
Author contributions
NYL.: Conceptualization, Formal analysis, Software, Writing—original draft, Writing—Review & editing, Visualisation. LP: Data preparation, Writing—Review & editing, Validation. AMW: Conceptualisation, Writing—Review & editing DV: Conceptualisation, Methodology, Software, Writing—original draft, Writing—Review & editing, Funding acquisition.
Competing financial interests
The authors declare no competing financial interests.
S1 Across-sessions within-subject testing
In the across-sessions within-subject testing method, we evaluate whether an experimental effect (e.g., a brain–behaviour relationship) changes significantly across multiple sessions. The key idea is to treat each session as a separate block and test if the effect sizes differ by session, using a permutation test that preserves session structure. Hence, we avoid shuffling raw data across sessions and instead permute at the level of model coefficients of each session. This design choice is essential because a direct permutation of the independent ($D$) and dependent ($R$) variables across sessions is invalid. Randomly reassigning trial data across sessions would disrupt the correspondence between stimulus and response within each session, and thereby violating exchangeability assumptions necessary for permutation-based inference.
A naive permutation approach would involve shuffling trials across sessions. This, however, would lead to condition mismatches: each session has a predefined stimulus order, meaning that swapping trials would misalign $D$ with its corresponding response variables $R$. Letting $R$ encode a binary condition (e.g., 00 for an inanimate stimulus and $1$ for an animate stimulus), a given trial at index $i$ in session $s$ might originally have $R^{(s)}_{i}=0$, but after permutation, that same index in another session $s^{\prime}$ could correspond to $R^{(s^{\prime})}_{i}=1$. This incorrect reassignment destroys the link between trial conditions and responses.
A naive approach might involve shuffling trials across sessions. This, however, would lead to condition mismatches: each session has a predefined stimulus order, meaning that swapping trials would misalign $D$ with its corresponding response variables $R$. Let $D$ represent the neural data and $R$ the behavioural conditions, with each trial indexed by $i$ within session $s$:
$D_{s,i}\sim R_{s,i}$ (1)
where $R_{s,i}\in{0,1}$ encodes the stimulus type (e.g., animate vs. inanimate). If $R$ is permuted across sessions, a given index $i$ in session $s$ that originally corresponded to $R_{s,i}=0$ may, after permutation, be reassigned to $R_{s^{\prime},i}=1$ in a different session $s^{\prime}$. This misalignment violates the experimental design, as trial positions across sessions no longer correspond to the same stimulus categories.
To illustrate, consider three sessions, each with four trials:
$D={D_{1,1},D_{1,2},D_{1,3},D_{1,4},D_{2,1},\dots,D_{3,4}}$ (2)
$R={1,0,0,1,1,1,0,0,0,1,0,1}$ (3)
If trials were randomly shuffled across sessions, the values in $R$ would be reassigned without considering the original session constraints. This destroys the association between $D$ and $R$ and disrupts the experimental structure.
Moreover, permuting entire sessions as indivisible blocks introduces additional inconsistencies. Each session has a fixed stimulus order, so if all trials from session $s$ were assigned to a different session $s^{\prime}$, stimulus labels and neural responses would no longer correspond to the same experimental events. This creates systematic errors, as responses originally evoked by one condition may be falsely associated with another condition. Furthermore, session lengths often vary, meaning the number of trials $N_{s}$ in session $s$ may not match that of another session $s^{\prime}$:
$N_{1}\neq N_{2}\neq\dots\neq N_{S}.$ (4)
When a shorter session is reassigned to a longer session’s, it results in an incomplete mapping, as some trials from the shorter session cannot fit within the longer session’s structure. Conversely, if a longer session’s data is assigned to a shorter session, its trials may inadvertently extend into another session’s range. This leads to unintended overlap, where trials from two originally separate sessions could be falsely merged in the permuted dataset, further distorting the null distribution.
\@fb@secFB
S1.1 Permutation Strategy: Beta Coefficients
To overcome these issues, we permute at the level of regression coefficients ($\beta$) rather than raw data. In practice, this means we first estimate session-specific $\beta$ weights that capture the $D$$R$ relationship within each session, then shuffle their assignments across sessions in a controlled manner for hypothesis testing. Formally, for each session $s$, we define $D^{(s)}$ (independent variables, e.g., brain data) and $R^{(s)}$ (dependent variables, e.g., experimental conditions or behavioural outputs). We obtain $\beta^{(s)}$ by fitting a regression model to that session’s data.
In practice, this means we first estimate session-specific $\beta$ weights that capture the $D$$R$ relationship within each session, then shuffle their assignments across sessions in a controlled manner for hypothesis testing. Formally, for each session $s$, we define $D^{(s)}$ (independent variables, e.g.,brain data) and $R^{(s)}$ (dependent variables, e.g., behavioural outputs). We obtain $\beta^{(s)}$ by fitting a regression model to a subset of that session’s data.
To ensure that the estimated $\beta$ coefficients generalise beyond the specific trials they were computed from, we split each session’s data into two parts:
  • A train set, used to estimate the session-specific $\beta$ coefficients.
  • A test set, used to compute a test statistic based on these estimated $\beta$ coefficients.
Using ordinary least squares on the train set, the session-specific coefficients are estimated as:
$\hat{\beta}^{(s)}=\big{(}D_{\text{train}}^{(s)\top}D_{\text{train}}^{(s)}\big{)}^{-1}D_{\text{train}}^{(s)\top}R_{\text{train}}^{(s)}.$ (5)
where $D^{(s)}{\text{test}}$ and $R^{(s)}{\text{test}}$ denote the held-out test data. We do not evaluate model performance; instead, we use the predicted responses $\hat{R}^{(s)}_{\text{test}}$ to compute an F-statistic:
$F^{(s)}=\frac{\text{Var}(\hat{R}^{(s)}{\text{test}})}{\text{Var}(R^{(s)}_{\text{test}}-\hat{R}^{(s)}_{\text{test}})}.$ (6)
These estimated coefficients are then applied to the test set for that session to compute predicted responses:
$\hat{R}_{\text{test}}^{(s)}=D_{\text{test}}^{(s)}\;\hat{\beta}^{(s)}.$ (7)
where $D_{\text{test}}^{(s)}$ and $R_{\text{test}}^{(s)}$ denote the held-out test data. We do not evaluate model performance; instead, we use the predicted responses $\hat{R}_{\text{test}}^{(s)}$ to compute an F-statistic, which serves as our test statistic. The F-statistic is based on the explained variance ($R^{2}$) between the predicted $\hat{R}_{\text{test}}^{(s)}$ and the actual $R_{\text{test}}^{(s)}$.The F-statistic, which quantifies how well the estimated model explains variance in the test set, is computed as:
Since we are evaluating the entire dataset as a whole, we compute the explained variance ($R^{2}$) and F-statistic using all test data from all sessions, rather than calculating separate statistics per session.
$\hat{R}_{\text{test}}=\big{[}\hat{R}_{\text{test}}^{(1)},\hat{R}_{\text{test}}^{(2)},\dots,\hat{R}_{\text{test}}^{(S)}\big{]}$ (8)
where $S$ is the total number of sessions. The explained variance ($R^{2}$) is then computed across the entire concatenated dataset:
$R^{2}=1-\frac{\sum(\hat{R}_{\text{test}}-R_{\text{test}})^{2}}{\sum(R_{\text{test}}-\bar{R}_{\text{test}})^{2}}.$ (9)
Finally, the F-statistic, which quantifies the proportion of variance explained by the model relative to residual variance, is computed as:
$F=\frac{(R^{2}/p)}{(1-R^{2})/(n-p)}$ (10)
where $p$ is the number of predictors, and $n$ is the total number of samples across all test sets.
\@fb@secFB
S1.2 Permutation Implementation
For our permutation test, we construct a null distribution by shuffling the assignment of regression coefficients ($\beta$) across sessions while maintaining session boundaries. In each permutation $\pi$, we randomly reassign the estimated coefficients $\hat{\beta}^{(s)}$ to a different session, generating a permuted coefficient $\hat{\beta}_{\pi}^{(s)}$. The corresponding predicted test responses for each session are then computed as:
$\hat{R}_{\text{test},\pi}^{(s)}=D_{\text{test}}^{(s)}\hat{\beta}_{\pi}^{(s)}.$ (11)
We aggregate these predictions across all sessions:
$\hat{R}_{\text{test},\pi}=\big{[}\hat{R}_{\text{test},\pi}^{(1)},\hat{R}_{\text{test},\pi}^{(2)},\dots,\hat{R}_{\text{test},\pi}^{(S)}\big{]}.$ (12)
From these predictions based on the permuted beta coefficients, we compute a new explained variance $R^{2}_{\pi}$ and F-statistic under the null hypothesis:
$R_{\pi}^{2}=1-\frac{\sum(\hat{R}_{\text{test},\pi}-R_{\text{test}})^{2}}{\sum(R_{\text{test}}-\bar{R}_{\text{test}})^{2}}.$ (13)
$F_{\pi}=\frac{(R^{2}_{\pi}/p)}{(1-R^{2}_{\pi})/(n-p)}.$ (14)
This process is repeated over multiple permutations, generating a distribution of null F-statistics $\{F_{\pi}\}$. The final p-value is computed as the proportion of permutations where the permuted statistic is greater than or equal to the observed statistic:
$p=\frac{\sum_{\pi=1}^{N_{\text{perm}}}\big{(}F_{\pi}\geq F_{0}\big{)}}{N_{\text{perm}}+1}.$ (15)
Here, $N_{\text{perm}}$ is the total number of permutations, and $F_{0}$ is the observed F-statistic (i.e., the first entry in the permutation distribution).
S2 Across-visits testing - Structured Monte Carlo resampling approach
The across-visits test assesses the relationship between a sequence of brain states, as identified by a Viterbi path, and a concurrent physiological or behavioural signal recorded simultaneously (e.g., pupil size, skin conductance, heart rate). Traditional permutation-based approaches, which shuffle time points or state assignments, fail in this context because they disrupt the temporal dependencies in the data, thereby violating exchangeability assumptions. Instead, we use a structured Monte Carlo resampling approach that preserves the transition structure of state sequences while generating a valid non-parametric test of statistical significance.
A key motivation for using this approach is that brain state sequences exhibit temporal dependencies. A naive shuffling of individual state labels or full time series would break the sequential order of state transitions, and thereby artificially distort the statistical properties of the data and leading to invalid inference. Unlike standard standard Monte Carlo simulations, which assume fixed transition probabilities, our framework requires dynapmically updated transitions. Therefore, we use a structured randomisation method that maintains the temporal properties of the original sequences while generating a valid null distribution.
\@fb@secFB
S2.1 Structured Monte Carlo Resampling Procedure
Rather than randomly permuting time points, we generate surrogate sequences by resampling entire state visit sequences while preserving their transition structure. This ensures that surrogate sequences retain realistic temporal properties. The procedure follows three key steps:
  1. 1.
    Preserve transition structure: Instead of assigning new states randomly at each time point, we maintain the original transition timing and only modify the sequence of state visits.
  2. 2.
    Exclude self-transitions: At each transition, a new state is selected randomly from the set of all states except for the previously occupied state. This prevents inflated self-transition probabilities.
  3. 3.
    Generate surrogate sequences: By repeating this process iteratively we generate a set of surrogate Viterbi paths that preserve the statistical properties of the original data while breaking the relationship between brain states and the external signal so that observed effects can be tested against a baseline where no real association exists. .
\@fb@secFB
S2.2 Example: Surrogate Viterbi Path Generation
To illustrate, consider a dataset where the original Viterbi path consists of three states (1, 2, and 3) recorded over 12 time points:
Observed Viterbi Path (Original Sequence)
$D=\{1,1,1,2,2,3,3,3,1,1,2,2\}$ (16)
For surrogate generation, we preserve the transition points (i.e., time points where the state changes) but randomly assign a new state at each transition. Hence, the previous state is not repeated:
Generated Surrogate Viterbi Path
$D_{\pi}=\{3,3,3,1,1,2,2,2,3,3,1,1\}$ (17)
Crucially, the transition structure remains unchanged, meaning:
  • The number of state visits per sequence remains identical.
  • State durations are preserved.
  • Transition timing is unchanged.
  • The mapping between $D$ and the external signal is disrupted, creating a valid null distribution.
This structured randomisation approach is designed to preserve the structure of each individual Viterbi path. Importantly, we did not shuffle individual time points or randomly reassign states at each time point (which would destroy temporal dependencies); instead, we only permuted the labels at the visit level, treating each entire sequence as a unit. Thus, each surrogate sequence is identical (as an ordered sequence of states) to some originally observed sequence from the data—it has the same segments, the same state transition points, and the same durations in each state as that real sequence. The only thing that has changed is which visit that sequence is associated with in the surrogate dataset. This ensures that state durations and transition patterns within each sequence are held constant across the randomisations, satisfying the need for structured (dependency-aware) resampling. This forms the basis for a non-parametric hypothesis test.
By generating multiple surrogate sequences, we construct a null distribution of the test statistic. This allows us to assess whether the observed association between brain states and the external physiological signal is stronger than expected by chance. The statistical significance of the observed effect is then computed by comparing it to this null distribution.